It it takes one minute for a passenger standing on an escalator to reach the top if the escalator does not move it takes him 3mintue to walk up how long will it take for the passenger to arrive at the top if he walk
 
 
up the moving escalator

Grade:11

1 Answers

Arun
5 years ago
Dear Vivek
 
Let the distance to the top be d.therefore, the speed of the man with escalator not moving = (d/1) m/min.therefore the speed of the escalator with the man at rest = (d/3) m/minWhen both the man and the escalator are moving, effective speed= (d/3 + d) m/min =(4d/3)m/minTherefore time taken when both man and escalator are moving = distance / speed = d/(4d/3) min. = 3/4 min = 45 seconds
 
Regards
Arun (askIITians forum expert)